# Deep Learning for Melt Pool Depth Contour Prediction From Surface Thermal Images via Vision Transformers

### v2

@@ Conclusion
In this work, a method for predicting the melt pool cross-sectional morphology from in-situ high-speed two-color thermal imaging of the melt pool surface was created. This is accomplished through the implementation of a hybrid CNN-Transformer model. By decomposing the model architecture into the spatial and temporal components of the prediction task, we leverage the advantages of both convolutional and attention-based architectures. Specifically, the stricter inductive bias of convolutional neural networks are leveraged to extract the spatial information contained within the individual melt pool frames via a ResNet backbone. Next, the temporal relationships between these encoded thermal image frames are learned by a transformer model to discover any potential temporal dependency within the image sequence.
The performance of this framework is evaluated based on the dimensions of the predicted melt pool, as well as shape-dependent metrics such as the Intersection-over-Union score. The optimal monitoring configuration for model performance is examined by benchmarking these metrics for both monochrome images and temperature estimates at different exposure times. Further comparisons are made between the performance of conventional computer vision architectures on this task, with our proposed model demonstrating greater stability than a U-Net or Vision Transformer-based architecture. Finally, we evaluate the utility of transfer-learning to reduce the amount of experimental data required for robust model training. The successful development of this model enables the prediction of the melt pool cross-section directly from in-situ monitoring information, allowing L-PBF practitioners to have robust early detection of lack-of-fusion and porosity estimates. Future extensions of this work could examine the temporal information that is learned by a sequential model at higher sampling rates for greater insight into the physics underlying the temporal oscillation of the melt pool. Additionally, experiments with powder applied to the build plate can yield additional insight into the variability of the observed melt pool during the build process.
@@ Appendix B FLOW-3D Simulation Details
We simulate a single-track bare plate melting simulation for each unique power-velocity combination in the FLOW-3D dataset. The simulated build plate is $1000\mu m\times 600\mu m\times 400\mu m$, in the $x$, $y$, and $z$ directions respectively. The $x$-axis indicates the laser direction of travel, while the $y$-axis lies along the width of the generated melt pool, and the $z$-axis is defined as the vertical axis. Each simulation is carried out for 500 $\mu s$ and data is reported at 5 $\mu s$ intervals, for a total of 100 time-dependent snapshots for each power-velocity combination. The material parameters used for simulation are reproduced in Table 3.
Table 3: Material Parameters used to simulate the SS316L melting process.
@@
Appendix C Experimental Data Variability
@@
Appendix D Machine Learning Architecture Details

### v3

@@ Conclusion
In this work, a method for predicting the melt pool cross-sectional morphology from in-situ high-speed two-color thermal imaging of the melt pool surface was created. This is accomplished through the implementation of a hybrid CNN-Transformer model. By decomposing the model architecture into the spatial and temporal components of the prediction task, we leverage the advantages of both convolutional and attention-based architectures. The stricter inductive bias of CNNs is used to extract spatial information from individual melt pool frames using a ResNet backbone. Next, a transformer model learned the temporal relationships between these encoded thermal image frames to discover temporal dependencies. Comparisons are made between the performance of architectures that do not consider temporal information, with our proposed model demonstrating greater stability, as shown in Figure 5.
The performance of this framework is evaluated based on the dimensions of the predicted melt pool, as well as shape-dependent metrics such as the Intersection-over-Union score. The implemented model is able to predict the contour within an average deviation of 10 $\mu m$, and a maximum deviation of 20 $\mu m$. By comparing the effect of the optical monitoring configuration, we note that the use of ratiometric temperature estimates increases the IoU over monochromatic images from 0.73 to 0.77. Finally, we evaluate the utility of transfer learning to reduce the amount of experimental data required for robust model training. By applying this strategy, predictions within an $R^{2}$ of 0.87 are achieved with a 50 % reduction in the number of annotated cross-sections necessary for analysis. The development of this model enables the prediction of the melt pool cross-section directly from in-situ monitoring information, allowing L-PBF practitioners to have robust early detection of lack-of-fusion and porosity estimates. Future extensions of this work could examine the temporal information that is learned by a sequential model at higher sampling rates for greater insight into the physics underlying the temporal oscillation of the melt pool. Additionally, experiments with powder applied to the build plate can yield additional insight into the variability of the observed melt pool during the build process.
@@ Appendix B FLOW-3D Simulation Details
We simulate a single-track bare plate melting simulation for each unique power-velocity combination in the FLOW-3D dataset. The simulated build plate is $1000\mu m\times 600\mu m\times 400\mu m$, in the $x$, $y$, and $z$ directions respectively. The $x$-axis indicates the laser direction of travel, while the $y$-axis lies along the width of the generated melt pool, and the $z$-axis is defined as the vertical axis. Each simulation is carried out for 500 $\mu s$ and data is reported at 5 $\mu s$ intervals, for a total of 100 time-dependent snapshots for each power-velocity combination. The power-velocity combinations used are identical to those reported in [30] for the SS316L dataset. The material parameters used for simulation are reproduced in Table 3.
Table 3: Material Parameters used to simulate the SS316L melt pool with the FLOW-3D model.
@@
Appendix C Eagar-Tsai Simulation Details
We also simulate a single-track bare plate melting simulation for each unique power-velocity combination in the Eagar-Tsai dataset. The simulated build plate is $1000\mu m\times 600\mu m\times 400\mu m$, in the $x$, $y$, and $z$ directions respectively. The simulation is solved to simulate the laser moving for 200 $\mu s$ at the specified velocity. The power of the laser is defined to vary from 50 W to 450 W at increments of 400 W. The velocity of the laser is defined to vary from 0.1 m/s to 1.55 m/s in increments of 0.05 m/s. The material parameters used for simulation are reproduced in Table 4.
Table 4: Material Parameters used to simulate the SS316L melt pool with the Eagar-Tsai model.
Parameter Value Units
Density, $\rho$ 8030 kg/m3
Thermal Conductivity, $k$ 16.2 W/m/K
Melting Point, $T_{m}$ 1658 K
Beam Radius, $\sigma$ 50 $\mu m$
Absorptivity, $A$ 0.3 -
Heat Capacity, $c_{p}$ 500 - J/kg K
Appendix D Experimental Data Variability
@@
Appendix E Machine Learning Architecture Details